You borrowed $70,000 in student loans. You plan to make monthly payments to repay the debt. The interest rate is fixed at 3.3% APR (with monthly compounding). a) If the loans are for 10 years, find the monthly payment. b) Suppose that you decide to pay $300 more per month instead of the required monthly payment. How long will it take to pay off the loan?
Answer #1

The monthly payments are the annuity amount on the loan . The loan amount is the present value of annuity.

Present value of aanuity = monthly payment * [ 1 - ( 1 + periodic int rate ) ^- no of periods ] / periodic int rate

70000 = MP * [ 1 - ( 1 + 0.033/12) ^-10*12 ] / 0.033/12

70000 = MP * 102.0911

MP= 70000 /102.0911 = 685.66

Monthly payment = 685.66

b) revised monthly payment = 685.66 + 300 = 985.66

70000 = 985.66 [ 1 - 1.033/12^- N ] / 0.033/12

Let us use Financial Calculato

PV= +- 70000

PMT = 985.66

I/Y = 0.275

N = ?

No of months = 79.12